Sam is 8 years older than his brother. Seven years ago, Sam was twice as old as his brother. How old is sam now?
Answer by macston(5194)

S=Sam's age; B=brother's age
.
S=B+8yrs Use this to substitute for S below.
.
S-7yrs=2(B-7yrs) Substitute for S.
B+8yrs-7yrs=2B-14yrs Subtract B from each side.
1yr=B-14yrs Add 14 years to each side.
15 years=B Brother is 15 years old.
.
S=B+8yrs
S=15yrs+8yrs
S=23yrs Sam is now 23 years old.
.
CHECK:
Seven years ago, Sam was twice as old as his brother.
S-7yrs=2(b-7yrs)
23yrs-7yrs=2(15yrs-7yrs)
16yrs=2(8yrs)
16yrs=16yrs